OpenStudy (ac3):

Change the cartesian integral into an equivalent polar integral. Then evaluate the polar integral.

OpenStudy (ac3):

\[\int\limits_{0}^{1}\int\limits_{0}^{\sqrt{1-Y ^{2}}}x ^{2}+y ^{2}dxdy\]

OpenStudy (ac3):

answer is pi/8

OpenStudy (ac3):

i got pi/4

OpenStudy (ikram002p):

oh show ur work then if possible :P

OpenStudy (ac3):

I believe the graph for sqrt(1-y^2) looks like this...

OpenStudy (ac3):



OpenStudy (ac3):

therefore the bounds for theta should be -pi/2 to pi/2

OpenStudy (ac3):

the circle has a radius of one so the bounds for r i say are 0-1

OpenStudy (ac3):

and then x^(2)+y^(2) is equal to r^2. so that changes to that

OpenStudy (ac3):

and so my double integral is...

OpenStudy (ac3):

\[\int\limits\limits_{-\pi/2}^{\pi/2}\int\limits\limits_{0}^{1}r ^{3}drd \theta\]

OpenStudy (ac3):

and it's r^3 because the formula says there is always an extra r inside the double integral when you switch from cartesian to polar

ganeshie8 (ganeshie8):

bounds for \(\theta\) are wrong

OpenStudy (ac3):

what are they supposed to be?

ganeshie8 (ganeshie8):

because you're misinterpreting the region of integration

ganeshie8 (ganeshie8):

the region is not a semicircle

OpenStudy (ac3):

it is that is the equation for a semi-circle

ganeshie8 (ganeshie8):